OpenStudy (anonymous):

Factor: \[\frac{ 2p-7 }{ 9 }^{3}\div \frac{ 7-2p }{ 3 }^{4}\]

OpenStudy (anonymous):

\[=\frac{ 2p-7^3 }{ 9 } \times \frac{ 3 }{ 7-2p^4 }=\frac{ 2p-7^3 }{ 3(7-2p^4 )}\]

OpenStudy (anonymous):

but you have to square all of the numbers...

OpenStudy (anonymous):

@genius12